# Random Utility with Aggregation

### v2

@@ 4.2 Condition on $\lambda$: menu-independence
The second condition that we discuss is that the distribution of aggregated categories is the same across choice sets, or markets.
@@
Given an aggregation function $X$, a composition distribution $\lambda$ is said to be menu independent if there exists $\lambda\in\Delta\Big(\prod_{a\in\mathscr{A}}(2^{X(a)}\setminus\{\emptyset\})\Big)$ such that for any $B\subseteq{\mathscr{A}}$ and $(S_{a})_{a\in B}\in\prod_{a\in B}(2^{X(a)}\setminus\{\emptyset\})$,
$\lambda_{B}((S_{a})_{a\in B})=\lambda\Big(\prod_{a\in B}(S_{a})\times\prod_{a\in\mathscr{A}\setminus B}(2^{X(a)}\setminus\{\emptyset\})\Big).$ (7)
@@
  • To explain the meaning of our condition, consider a stronger condition as follows: there exists $\lambda^{a}\in\Delta(2^{X(a)}\setminus\{\emptyset\})$ for each $a\in{\mathscr{A}}$ such that for any $B\subseteq{\mathscr{A}}$ and $(S_{a})_{a\in B}\in\prod_{a\in B}(2^{X(a)}\setminus\{\emptyset\})$,
    $\lambda_{B}((S_{a})_{a\in B})=\prod_{a\in B}\lambda^{a}(S_{a}).$ (8)
    The condition (7) says that there exists a joint distribution that works across choice sets. [^15] For example, consider $a,b\in\mathscr{A}_{A}$ and let $X(a)=\{x_{1},y_{1}\}$ and $X(b)=\{x_{2},y_{2}\}$. Definition 2.5 allows the situation that $a$ means alternative $x_{1}$ ($y_{1}$) if and only if $b$ means alternative $x_{2}$ ($y_{2}$, respectively); and $a$ means $x_{1}$ and $y_{1}$ equally likely. In this example, the probability that $a$ means $x_{1}$ and $b$ means $y_{2}$ is zero, which is not allowed in (8). (7) is weaker than condition (8) in the sense that it allows correlation across aggregated categories. The two conditions coincide when there is a single aggregated category. In that case, both reduce to menu-independent aggregation: $\lambda_{D}=\lambda_{E}$ for all $D,E\subseteq\mathscr{A}$.
  • By the finiteness and the additivity, for any $B\subseteq{\mathscr{A}}$ and $(S_{a})_{a\in B}\in\prod_{a\in B}(2^{X(a)}\setminus\{\emptyset\})$, the menu independence implies
    $\lambda_{B}((S_{a})_{a\in B})=\sum_{(S_{c})_{a\in B^{c}}\in\prod_{c\in B^{c}}(2^{X(c)}\setminus\{\emptyset\})}\lambda((S_{a})_{a\in B},(S_{c})_{c\in B^{c}}).$
@@
Proposition 4.5 shows that if the composition distribution $\lambda$ is menu independent, then each aggregated category can be treated as a single good. Whether this assumption holds depends on the setting. In some markets, $\lambda$ may indeed be menu independent. A natural example is automobiles: when alternatives are aggregated by brand, the distribution of models within a brand is relatively stable, conditional on that brand being available, and thus invariant to which other brands are offered. In other markets, however, $\lambda$ is menu dependent. As discussed, the outside option often varies across markets (e.g., with income, store access, or local amenities), so its composition—and hence $\lambda_{A}$—changes with the menu $A$.
@@ 5 Simulation
These results motivate the following conjecture: As the composition distribution $\lambda$ departs from menu independence, or as preferences depart from the non-overlapping condition, the likelihood that $\rho$ can be rationalized by an ARUM diminishes. Consequently, estimation based on an ARUM, such as the logit model, is expected to yield increasingly biased utility estimates. This section empirically investigates this conjecture through simulation analysis and also examines how large such biases can be, an important practical question.
Specifically, we generate true unobservable choice data $\rho_{\scriptscriptstyle\mathscr{X}}$ defined over the set $\mathscr{X}$ of underlying alternatives using a logit model with fixed utilities. Given $\rho_{\scriptscriptstyle\mathscr{X}}$ and a specified composition distribution $\lambda$ as well as an aggregation function $X$, we construct the observed dataset $\rho$ on the set $\mathscr{A}$ of categories. We then estimate utility parameters from $\rho$ using a logit specification and measure the resulting estimation bias. In addition, we compute the distance from $\rho$ to the set of ARUMs (i.e., the ARU polytope) to evaluate the extent of rationality loss.
Our analysis is organized around the following conjectures:
  1. (A)
    As $\lambda$ becomes more menu dependent or preferences more overlapping, $\rho$ is increasingly unlikely to be rationalized by an ARUM, let alone by a logit model defined over $\mathscr{A}$. This leads to larger estimation biases (see Section 5.2).
  2. (B)
    The reason for Conjecture (A) is that the distance from $\rho$ to the ARU polytope grows with greater menu dependence or preference overlap (see Section 5.3).
We first confirm Conjecture (A). In particular, we demonstrate that the resulting biases are often substantial enough not only to distort estimated utility levels but also to overturn inferred preference orderings. We then demonstrate that Conjecture (B) explains why this occurs.
To establish these conjectures, we fix certain parameter values that are not the main focus of the analysis. Section B of the online appendix reports simulations using alternative parameter values.
@@ 5.1 Setup
Consider three alternatives $x,y,a_{0}$, where $x,y$ are non-aggregated categories [^16] For simplcity, we abuse the notation $x$ and $y$ to denote both the non-aggregated categories and their respective underlying alternatives.; $a_{0}$ is an aggregated category (the outside option). The outside option can mean good $z$ or $w$. That is, $X(a_{0})=\{z,w\}$. Given fixed utility values $u(x),u(y),u(z)$, and $u(w)$, we generate the true dataset $\rho_{\scriptscriptstyle\mathscr{X}}$ over the set $\mathscr{X}$ of underlying alternatives by assuming the logit model: for all $D\subseteq\{x,y,z,w\}$ and $i\in D$,
$\rho_{\scriptscriptstyle\mathscr{X}}(D,i)=\frac{\exp(u(i))}{\sum_{j\in D}\exp(u(j))}.$
We examine three distinct markets, each represented by one of the following choice sets: $\{x,a_{0}\}$, $\{y,a_{0}\}$, and $\{x,y,a_{0}\}$. Importantly, the composition of the aggregated category $a_{0}$ varies across these markets; that is, $\lambda_{\{x,a_{0}\}}$, $\lambda_{\{y,a_{0}\}}$, and $\lambda_{\{x,y,a_{0}\}}$ over $\{\{w\},\{z\},\{w,z\}\}$ are allowed to differ. With $\rho_{\scriptscriptstyle\mathscr{X}}$ and $\lambda$, we construct the observable reduced dataset $\rho$ over the set $\mathscr{A}$ of aggregated categories. That is for any $D\subseteq\{x,y\}$ and $i\in D$, we define $\rho(D\cup a_{0},i)$ as
@@
Given the observed aggregated dataset $\rho$, we estimate the utility parameters $\hat{u}(x)$ and $\hat{u}(y)$ under the assumption that $\rho$ follows a multinomial logit model defined over $\mathscr{A}$ and that the utility of the outside option is zero, i.e., $\exp(u(a_{0}))=1$. Note that in this setup
  • The composition distribution $\lambda$ becomes more menu dependent as $\lambda_{\{x,a_{0}\}}$, $\lambda_{\{y,a_{0}\}}$, and $\lambda_{\{x,y,a_{0}\}}$ become more distinct.
  • Fixing the true utilities of $u(x)$ and u(y)u(y)—say, u(x)>u(y)u(x)>u(y)—the preference structure becomes more overlapping as $\max\{u(z),u(w)\}$ exceeds $u(x)$ and $\min\{u(z),u(w)\}$ falls below $u(y)$.
@@ 5.2 Estimation Bias from ARUM Misspecification
To confirm Conjecture (A), we quantify the estimation bias using the following measure:
@@
where $\hat{u}(x)$ and $\hat{u}(y)$ denote the maximum-likelihood estimates of $u(x)$ and $u(y)$, respectively. This measure captures the distortion in relative utilities between alternatives $x$ and $y$. Remember that in our setup the values of $u(x)$ and $u(y)$ are fixed say $u(x)=2$ and $u(y)=1$. We are interested in the values of the measure as well as the sign of the measure. When the measure is less than 1-1 it means that in the estimate we have $\hat{u}(y)>\hat{u}(x)$, even though in the true data we have $u(x)>u(y)$. This means that bias affects not only the estimated utility levels but also the inferred preference orderings.
The values of the measure are also meaningful. To see this, consider the case in which $\hat{u}(x)-\hat{u}(y)>u(x)-u(y)$. Then, by taking the exponential of the measure, we have
$\frac{\exp(\hat{u}(x))}{\exp(\hat{u}(y))}\Big/\frac{\exp(u(x))}{\exp(u(y))},$
which represents the ratio of the estimated odds ratio to the true odds ratio of choosing $x$ over $y$. A deviation in this ratio reflects how the estimated utilities—and thus predicted behavior—diverges from the true underlying model.
@@ 5.2.1 Effect of composition distribution $\lambda$ on bias
In the first simulation, we test Conjecture (A) focusing on the role of $\lambda$: the bias defined in (9) becomes larger as $\lambda$ becomes more menu dependent (i.e., as $\lambda$ changes across choice sets $\{x,a_{0}\},\{y,a_{0}\}$, $\{x,y,a_{0}\}$). As the preference structure is not the focus of the analysis, we fixed the values $u(x)=2$, $u(y)=1$, $u(z)=3,u(w)=0$.
@@
Figure 1: Heatmap of biases across values of $\lambda_{\{x,a_{0}\}}$, with $\lambda_{\{y,a_{0}\}}$ and $\lambda_{\{x,y,a_{0}\}}$ fixed at $(0.8,0.1,0.1)$ and utilities fixed at $u(x)=2$, $u(y)=1$, $u(z)=3$, $u(w)=0$. The horizontal axis reports $\lambda_{\{x,a_{0}\}}(\{z\})$, the vertical axis reports $\lambda_{\{x,a_{0}\}}(\{w\})$, and color intensity represents the magnitude of the bias defined in (9). The blue-outlined cell marks the independent case, where $\lambda_{\{x,a_{0}\}}=\lambda_{\{y,a_{0}\}}=\lambda_{\{x,y,a_{0}\}}$; cells farther from the blue cell correspond to more menu-dependent cases, where the bias is larger.
@@
First, we construct a heatmap to visualize the sensitivity of bias to deviations from the independence case. In Figure 1, we vary the values of $\lambda_{\{x,a_{0}\}}$ while equating and holding $\lambda_{\{x,y,a_{0}\}}$ and $\lambda_{\{y,a_{0}\}}$ fixed at (0.8, 0.1, 0.1)(0.8,\ 0.1,\ 0.1), where the first, second, and third entries represent the probabilities assigned to $\{z\}$, $\{w\}$, and $\{z,w\}$, respectively. Note that the horizontal axis measures $\lambda_{\{x,a_{0}\}}(\{z\})$, while the vertical axis measures $\lambda_{\{x,a_{0}\}}(\{w\})$. (Note that the numbers $\lambda_{\{x,a_{0}\}}(\{z\})$ and $\lambda_{\{x,a_{0}\}}(\{w\})$ determine $\lambda_{\{x,a_{0}\}}$ uniquely as $\lambda_{\{x,a_{0}\}}(\{z,w\})=1-\lambda_{\{x,a_{0}\}}(\{w\})-\lambda_{\{x,a_{0}\}}(\{z\})$.) Each grid point in the heatmap represents a specific assignment of $\lambda_{\{x,a_{0}\}}$; the bias is computed at each grid point. The blue-outlined cell corresponds to the independent case, in which $\lambda_{\{x,a_{0}\}}(\{z\})=0.8$ and $\lambda_{\{x,a_{0}\}}(\{w\})=0.1$, implying that $\lambda_{\{x,a_{0}\}}=\lambda_{\{y,a_{0}\}}=\lambda_{\{x,y,a_{0}\}}$.
@@
The heatmap supports Conjecture (A): bias grows as $\lambda$ becomes more menu dependent—that is, as the cell moves further away from the independent case. The mechanism is straightforward: positive bias arises when $x$ is overvalued relative to $a_{0}$ in the menu $\{x,a_{0}\}$, while negative bias arises when $x$ is undervalued relative to $a_{0}$ in the menu $\{x,a_{0}\}$. When $a_{0}$ consists of $\{w\}$, it is relatively unattractive, whereas when it consists of $\{z\}$ or $\{z,w\}$, it is more attractive, since in our setup $u(z)$ is high and $u(w)$ is low. Hence, the large positive bias in the figure occurs when $a_{0}$ is composed of w{w} more often in the menu $\{x,a_{0}\}$ (e.g., when $\lambda_{\{x,a_{0}\}}(\{w\})=1$). Conversely, the negative bias arises when $a_{0}$ is composed of $\{z\}$ or $\{z,w\}$ (e.g., when $\lambda_{\{x,a_{0}\}}(\{w\})=0$).

### v3

@@ 4.2 Condition on $\lambda$: menu-independence
The second condition is on the composition distribution:
@@
Given an aggregation correspondence $X$, a composition distribution $(\lambda_{B})_{B\in\mathscr{D}}$ is said to be menu independent if there exists an unconditional distribution $\lambda\in\Delta\Big(\prod_{a\in\mathscr{A}}(2^{X(a)}\setminus\{\emptyset\})\Big)$ that the marginal distribution of $\lambda$ on $\prod_{a\in B}(2^{X(a)}\setminus\{\emptyset\}$ coincides with $\lambda_{B}$ for all $B\in\mathscr{A}$. [^19] When $(\lambda_{B})_{B\in\mathscr{D}}$ is menu-independent and when $\mathscr{A}\in\mathscr{D}$, the unconditional distribution must coincides with $\lambda_{\mathscr{A}}$. Thus the menu-independence condition can be written equivalently as the marginal distribution of $\lambda_{\mathscr{A}}$ on $\prod_{a\in B}(2^{X(a)}\setminus\{\emptyset\})$ coincides with $\lambda_{B}$ for all $B\in\mathscr{A}$.
@@
The menu-independence requires that the conditional distribution must be consistent across menus. If $(\lambda_{B})_{B\in\mathscr{D}}$ is menu-independent, then for any $A,B\in\mathscr{D}$, the marginal distributions $\lambda_{A}$ and $\lambda_{B}$ on $\prod_{a\in A\cap B}(2^{X(a)}\setminus\{\emptyset\})$ must coincide. In particular, $\lambda_{A}=\lambda_{B}$ whenever $A\cap\mathscr{A}_{N}=B\cap\mathscr{A}_{N}$.
On the other hand, the menu-independence allows correlation across aggregates. To see this, consider a stronger condition that does not allow the correlation: there exists $\lambda^{a}\in\Delta(2^{X(a)}\setminus\{\emptyset\})$ for each aggregate $a\in{\mathscr{A}}$ such that for any $B\subseteq{\mathscr{A}}$ and $(S_{a})_{a\in B}\in\prod_{a\in B}(2^{X(a)}\setminus\{\emptyset\})$, $\lambda_{B}((S_{a})_{a\in B})=\prod_{a\in B}\lambda^{a}(S_{a})$. [^20] For example, consider $a,b\in\mathscr{A}_{N}$ and let $X(a)=\{x_{1},y_{1}\}$ and $X(b)=\{x_{2},y_{2}\}$. Definition 4.2 allows the situation that $a$ means alternative $x_{1}$ ($y_{1}$) if and only if $b$ means alternative $x_{2}$ ($y_{2}$, respectively); and $a$ means $x_{1}$ and $y_{1}$ equally likely. In this example, the probability that $a$ means $x_{1}$ and $b$ means $y_{2}$ is zero, which is not allowed in the stronger menu-independence condition. The two conditions coincide when there is only one non-atomic aggregate. In that case, both reduce to the following: $\lambda_{D}=\lambda_{E}$ for all $D,E\subseteq\mathscr{A}$.
@@
Proposition 4.2 shows that, without any assumption on $\mu_{\mathscr{X}}$, if the composition distribution $\lambda$ is menu independent, then each aggregate can be treated as a single good. Whether this assumption holds depends on the setting. In some markets, $\lambda$ may indeed be menu independent. A natural example is automobiles: when alternatives are aggregated by brand, the distribution of models within a brand is relatively stable, conditional on that brand being available, and thus invariant to which other brands are offered. In other markets, however, $\lambda$ is menu dependent. As discussed, the outside option often varies across markets (e.g., with income, store access, or local amenities), so its composition—and hence $\lambda_{A}$—changes with the menu $A$.
4.3 How to define outside option?
The previous two propositions provide guidance on how to define the outside option. Specifically, an underlying alternative $x$ should be excluded from the outside option and instead modeled as a separate aggregate if either:
(1) the desirability of $x$ is very different from the other underlying alternatives that the outside option $a_{0}$ represents, since this would violate non-overlappingness (Proposition 4.1); or
(2) the availability of $x$ varies across markets, since including it in the outside option would violate menu-independence (Proposition 4.2).
@@ 5 Simulation
Our theoretical results suggest that if $\lambda$ is sufficiently menu dependent and $\mu_{\mathscr{X}}$ is sufficiently overlapping, $\rho$ is unlikely to be rationalized by an ARUM. In this section, we identify two empirical observations based on the implication of the theoretical results: if $\lambda$ is sufficiently menu dependent and $\mu_{\mathscr{X}}$ is sufficiently overlapping, then (a) estimation bias becomes large, and (b) the distance from the RU polytope becomes large.
We also examine how large such biases can be, an important practical question. In particular, we demonstrate that the resulting biases are often substantial enough not only to distort estimated utility levels but also to overturn inferred preference orderings.
@@ 5.1 Setup
Consider three alternatives $x,y,a_{0}$, where $x,y$ are atomic aggregates  [^21] For simplicity, we abuse the notation $x$ and $y$ to denote both the aggregates and their respective underlying alternatives.; $a_{0}$ is a non-atomic aggregate (the outside option). The outside option can mean good $z$ or $w$. That is, $X(a_{0})=\{z,w\}$. [^22] This means that the set of RU-rational stochastic choice functions is $RU(2)$. We adopt this setup for two reasons. First, it permits a two-dimensional graphical representation of our bias and distance results. Second, as shown in Theorem 3.3, the set of RU-rational stochastic choice functions is smallest in $RU(2)$, although it remains strictly larger than the ARU polytope. We show that even in this setting, the biases and distances are substantial. With a larger number of underlying alternatives, both the biases and the distances would increase. Given fixed utility values $u(x),u(y),u(z)$, and $u(w)$, we generate the true dataset $\rho_{\scriptscriptstyle\mathscr{X}}$ over the set $\mathscr{X}$ of underlying alternatives by assuming the logit model: for all $D\subseteq\{x,y,z,w\}$ and $i\in D$, $\rho_{\scriptscriptstyle\mathscr{X}}(D,i)=\exp(u(i))/\sum_{j\in D}\exp(u(j))$.
We examine three distinct markets, each represented by one of the following choice sets: $\{x,a_{0}\}$, $\{y,a_{0}\}$, and $\{x,y,a_{0}\}$. Importantly, the composition of the aggregate $a_{0}$ varies across these markets; that is, $\lambda_{\{x,a_{0}\}}$, $\lambda_{\{y,a_{0}\}}$, and $\lambda_{\{x,y,a_{0}\}}$ over $\{\{w\},\{z\},\{w,z\}\}$ are allowed to differ. With $\rho_{\scriptscriptstyle\mathscr{X}}$ and $\lambda$, we construct the observable reduced dataset $\rho$ over the set $\mathscr{A}$ of aggregates. That is for any $D\subseteq\{x,y\}$ and $i\in D$, we define $\rho(D\cup a_{0},i)$ as
@@
Given the observed aggregated dataset $\rho$, we estimate the utility parameters $\hat{u}(x)$ and $\hat{u}(y)$ under the assumption that $\rho$ follows a multinomial logit model defined over $\mathscr{A}$ and that the utility of the outside option is zero, i.e., $\exp(u(a_{0}))=1$. Note that in this setup, (i) the composition distribution $\lambda$ becomes more menu dependent as $\lambda_{\{x,a_{0}\}}$, $\lambda_{\{y,a_{0}\}}$, and $\lambda_{\{x,y,a_{0}\}}$ become more distinct; (ii) Fixing the true utilities of $u(x)$ and u(y)u(y)—say, u(x)>u(y)u(x)>u(y)—the preference structure becomes more overlapping as $\max\{u(z),u(w)\}$ exceeds $u(x)$ and $\min\{u(z),u(w)\}$ falls below $u(y)$.
In the following sections, we fix certain parameter values that are not the main focus of the analysis. Section E of the online appendix reports simulations under alternative parameter values, with qualitatively similar results throughout.
@@ 5.2 Estimation Bias from ARUM Misspecification
We quantify the estimation bias using the following measure:
@@
where $\hat{u}(x)$ and $\hat{u}(y)$ denote the maximum-likelihood estimates of $u(x)$ and $u(y)$, respectively. This measure captures the distortion in relative utilities between alternatives $x$ and $y$. Remember that in our setup the values of $u(x)$ and $u(y)$ are fixed at $u(x)=2$ and $u(y)=1$. We are interested in the values of the measure as well as the sign of the measure. When the bias (7) is less than 1-1 it means that in the estimate we have $\hat{u}(y)>\hat{u}(x)$, even though in the true data we have $u(x)>u(y)$. This means that bias affects not only the estimated utility levels but also the inferred preference orderings.
The values of the measure are also meaningful. To see this, consider the case in which $\hat{u}(x)-\hat{u}(y)>u(x)-u(y)$. Then, by taking the exponential of the measure, we have $\frac{\exp(\hat{u}(x))}{\exp(\hat{u}(y))}\Big/\frac{\exp(u(x))}{\exp(u(y))}$, which represents the ratio of the estimated odds ratio to the true odds ratio of choosing $x$ over $y$. A deviation in this ratio reflects how the estimated utilities—and thus predicted behavior—diverges from the true underlying model.
@@ 5.2.1 Effect of composition distribution $\lambda$ on bias
In the first simulation, focusing on the role of $\lambda$: the bias defined in (7) becomes larger as $\lambda$ becomes more menu dependent (i.e., as $\lambda$ changes across choice sets $\{x,a_{0}\},\{y,a_{0}\}$, $\{x,y,a_{0}\}$). As the preference structure is not the focus of the analysis, we fixed the values $u(x)=2$, $u(y)=1$, $u(z)=3,u(w)=0$.
@@
a Heatmap of biases across values of $\lambda_{\{x,a_{0}\}}$, with $\lambda_{\{y,a_{0}\}}$ and $\lambda_{\{x,y,a_{0}\}}$ fixed at $(0.8,0.1,0.1)$ and utilities fixed at $u(x)=2$, $u(y)=1$, $u(z)=3$, $u(w)=0$. The horizontal axis reports $\lambda_{\{x,a_{0}\}}(\{z\})$, the vertical axis reports $\lambda_{\{x,a_{0}\}}(\{w\})$, and color intensity represents the magnitude of the bias defined in (7). The blue-outlined cell marks the independent case, where $\lambda_{\{x,a_{0}\}}=\lambda_{\{y,a_{0}\}}=\lambda_{\{x,y,a_{0}\}}$; cells farther from the blue cell correspond to more menu-dependent cases, where the bias is larger.
b Heatmap of biases across values of $u(z)$ and $u(w)$, with $u(x)=2$, $u(y)=1$, and composition distributions fixed at $\lambda_{\{x,y,a_{0}\}}=(0.8,0.1,0.1)$, $\lambda_{\{x,a_{0}\}}=(0.8,0.1,0.1)$, and $\lambda_{\{y,a_{0}\}}=(0.1,0.8,0.1)$. The horizontal axis reports $u(z)$, the vertical axis reports $u(w)$, and color intensity indicates the magnitude of the bias defined in (7) (red for positive bias, blue for negative bias). Biases are largest when preferences overlap strongly, i.e., when $u(z)$ and $u(w)$ straddle $u(x)=2$ and $u(y)=1$.
@@
First, we construct a heatmap to visualize the sensitivity of bias to deviations from the independence case. In Figure 2a, we vary the values of $\lambda_{\{x,a_{0}\}}$ while equating and holding $\lambda_{\{x,y,a_{0}\}}$ and $\lambda_{\{y,a_{0}\}}$ fixed at (0.8, 0.1, 0.1)(0.8,\ 0.1,\ 0.1), where the first, second, and third entries represent the probabilities assigned to $\{z\}$, $\{w\}$, and $\{z,w\}$, respectively. Note that the horizontal axis measures $\lambda_{\{x,a_{0}\}}(\{z\})$, while the vertical axis measures $\lambda_{\{x,a_{0}\}}(\{w\})$. (Note that the numbers $\lambda_{\{x,a_{0}\}}(\{z\})$ and $\lambda_{\{x,a_{0}\}}(\{w\})$ determine $\lambda_{\{x,a_{0}\}}$ uniquely as $\lambda_{\{x,a_{0}\}}(\{z,w\})=1-\lambda_{\{x,a_{0}\}}(\{w\})-\lambda_{\{x,a_{0}\}}(\{z\})$.) Each grid point in the heatmap represents a specific assignment of $\lambda_{\{x,a_{0}\}}$; the bias is computed at each grid point. The blue-outlined cell corresponds to the menu-independent case, in which $\lambda_{\{x,a_{0}\}}(\{z\})=0.8$ and $\lambda_{\{x,a_{0}\}}(\{w\})=0.1$, implying that $\lambda_{\{x,a_{0}\}}=\lambda_{\{y,a_{0}\}}=\lambda_{\{x,y,a_{0}\}}$.
@@
The heatmap shows that bias grows as $\lambda$ becomes more menu dependent—that is, as the cell moves further away from the menu-independent case. The mechanism is straightforward: positive bias arises when $x$ is overvalued relative to $a_{0}$ in the menu $\{x,a_{0}\}$, while negative bias arises when $x$ is undervalued relative to $a_{0}$ in the menu $\{x,a_{0}\}$. When $a_{0}$ consists of $\{w\}$, it is relatively unattractive, whereas when it consists of $\{z\}$ or $\{z,w\}$, it is more attractive, since in our setup $u(z)$ is high and $u(w)$ is low. Hence, the large positive bias in the figure occurs when $a_{0}$ is composed of w{w} more often in the menu $\{x,a_{0}\}$ (e.g., when $\lambda_{\{x,a_{0}\}}(\{w\})=1$). Conversely, the negative bias arises when $a_{0}$ is composed of $\{z\}$ or $\{z,w\}$ (e.g., when $\lambda_{\{x,a_{0}\}}(\{w\})=0$).
5.2.2 Effect of preference structure on bias
In the last simulation on biases, we vary the preference structure to evaluate how overlapping preferences affect bias. We assume that the true unobservable dataset $\rho_{\scriptscriptstyle\mathscr{X}}$ is generated by logit models with $u(x)=2$, $u(y)=1$, and various values of $u(z)$ and $u(w)$ on a uniform grid over [5,5][-5,5]. The composition distributions $\lambda$ are fixed to be menu-dependent, since $\lambda$ is not the focus of this analysis. Specifically, we set $\lambda_{\{x,y,a_{0}\}}=(0.8,0.1,0.1)$, $\lambda_{\{x,a_{0}\}}=(0.8,0.1,0.1)$, and $\lambda_{\{y,a_{0}\}}=(0.1,0.8,0.1)$, where the three coordinates represent the probabilities of $\{z\}$, $\{w\}$, and $\{z,w\}$, respectively.
Figure 2b displays the resulting heatmap, with $u(z)$ and $u(w)$ on the axes. Red indicates positive bias at the top left corner and blue indicates negative bias at the bottom right corner, with color intensity reflecting the magnitude. The figure again confirms that bias is largest when the preference structure is highly overlapping. In particular, as $u(z)$ and $u(w)$ move farther apart, cells lie farther from the diagonal, and the magnitude of bias increases. Conversely, when both $u(z)$ and $u(w)$ are either very high or very low, the bias decreases. We provide further explanation in Section E.1 of the online appendix.